# Temporal Preference Optimization for Long-Form Video Understanding

### v2

@@ Abstract
Despite significant advancements in video large multimodal models (video-LMMs), achieving effective temporal grounding in long-form videos remains a challenge for existing models. To address this limitation, we propose Temporal Preference Optimization (TPO), a novel post-training framework designed to enhance the temporal grounding capabilities of video-LMMs through preference learning. TPO adopts a self-training approach that enables models to differentiate between well-grounded and less accurate temporal responses by leveraging curated preference datasets at two granularities: localized temporal grounding, which focuses on specific video segments, and comprehensive temporal grounding, which captures extended temporal dependencies across entire video sequences. By optimizing on these preference datasets, TPO significantly enhances temporal understanding while reducing reliance on manually annotated data. Extensive experiments on three long-form video understanding benchmarks—LongVideoBench, MLVU, and Video-MME—demonstrate the effectiveness of TPO across two state-of-the-art video-LMMs. Notably, LLaVA-Video-TPO establishes itself as the leading 7B model on the Video-MME benchmark, underscoring the potential of TPO as a scalable and efficient solution for advancing temporal reasoning in long-form video understanding. Project page: https://ruili33.github.io/tpo_website.
@@ 1 Introduction
Figure 1: Temporal Preference Optimization (TPO) is a self-improvement preference optimization technique designed to enhance video comprehension in video-LMMs by modeling temporal preferences at two granular levels: localized and comprehensive TPO. In localized TPO (upper-left), we generate queries focused on short segments, with contrastive responses that retain or exclude the target segment. For comprehensive TPO (lower-left), queries are designed for high-level understanding, using intact video versus sparse downsampled video for contrasting responses. After post-filtering, the contrast response pairs are serving as the preference dataset to train a video-LMM, guiding the model to prioritize preferred responses for improved video understanding.
@@
Recent advancements in video large multimodal models (video-LMMs) [52, 2, 45] have marked a significant step forward for generalizable video understanding. While image-based LMMs (image-LMMs) [20, 5, 39] primarily focus on spatial reasoning, video-LMMs face the additional complexity of modeling temporal dependencies—a critical aspect for capturing the dynamic nature of video content.
Current video-LMM approaches often follow a two-stage training paradigm: pre-training on large-scale multimodal datasets, followed by supervised fine-tuning (SFT) [6] on curated video-text instruction tuning datasets [7, 67]. This process typically involves dense captioning of video frames using cutting-edge large image-LMMs like GPT-4o [2]. These captions are then transformed into question-answer pairs by leveraging large language models (LLMs). The hope is that this two-stage training enables the model to align video content with LLM capabilities via the text interaction, ultimately learning instruction-following behavior for both general and temporal understanding tasks.
However, these methods for temporal modeling have significant challenges. First, they depend on large-scale, meticulously curated instruction-tuning datasets, which rely heavily on costly advanced models for dataset generation. Second, these approaches lack explicit optimization for temporal grounding capabilities, as the training signals primarily originate from weakly aligned video content and instruction-tuning pairs. As a result, these methods often fail to capture nuanced temporal relationships, particularly in long-form videos, leading to suboptimal performance on tasks requiring fine-grained or long-context temporal grounding. Zhang et al. [66] takes a step forward by utilizing Direct Preference Optimization (DPO) [44] to align the outputs of video-LMMs with those of advanced LLMs. However, this approach still evaluates preferences based solely on textual inputs, overlooking the complex temporal information inherent in the video modality. These limitations underscore the urgent need for more effective and efficient strategies to enhance temporal understanding, particularly for long-form video comprehension.
In this work, we introduce Temporal Preference Optimization (TPO), a self-training framework designed to enhance the temporal grounding capabilities of video-LMMs. Under the preference learning framework, TPO utilizes the contrast of the preferred and dis-preferred responses to encourage the model to prioritize preferred and temporally grounded responses. To achieve this, TPO injects two different granularities of temporal preference into the preference dataset: localized temporal grounding and comprehensive temporal grounding. For localized temporal grounding, questions target specific video segments, with preferred responses derived from the corresponding segment and dis-preferred responses from unrelated segments. Comprehensive temporal grounding, in contrast, involves broader video content, where preferred responses are generated using the full video, and dis-preferred responses from a subsampled version that omits key information. By manipulating the video input according to the input query, we are able to automatically inject desired temporal preference into the generated preference dataset via simple transformations of the input video. The resulting preference dataset is then used to optimize the video-LMM’s temporal grounding capabilities through Direct Preference Optimization (DPO) [44], a method chosen for its flexibility and stability. This approach ensures that TPO effectively enhances temporal reasoning across both fine-grained and long-context scenarios, making it a robust solution for advancing long-form video understanding tasks.
@@
We conducted extensive experiments on three challenging long-form video understanding benchmarks, and the results clearly demonstrate that TPO significantly improves the temporal understanding capabilities of video-LMMs, achieving the performance boost of 2.9%, 3.1%, and 2.5% on LongVideoBench [55], MLVU [70], and Video-MME [15] on top of state-of-the-art video-LMM LongVA-7B [65]. Furthermore, LLaVA-Video-TPO achieves the strongest 7B model on the Video-MME benchmark.
@@ 2 Preliminaries
Video large multimodal model (video-LMM) [52, 34, 32, 67] usually consists of a large language model (LLM) $\pi_{\theta}$, a visual encoder $f_{V}$ and a multimodal projector $f_{p}$. Video-LMM accepts the input of a video $V$ and a sequence of input text $\mathbf{x}$. The input video is first tokenized into the visual token $X_{V}$ by video-LMM’s visual encoder, which is then projected into the textual embedding space by the multimodal projector $\mathbf{v}=f_{p}(f_{V}(V))$. Such video-LMM models the probability of a response $\mathbf{y}$:
$P(\mathbf{y}|\mathbf{x},V)=\prod_{i=1}^{n}\pi_{\theta}(y_{i}|y_{<i},\mathbf{x},\mathbf{v})$ (1)
In our experiments, we implement TPO based on two widely-used video-LMMs, LongVA-7B [65] and LLaVA-Video-7B [68], leveraging their publicly available open-source checkpoints. LongVA-7B is specifically designed for long-form video comprehension by adapting methodologies from long-text processing, providing a robust foundation for our work. LLaVA-Video-7B is one of the open-source state-of-the-art video-LMMs.
Preference learning [41, 49, 72] focuses on modeling human preferences to align model behavior with user expectations. In LLMs and image-LMMs, this involves training models to generate responses favored by users. This is typically achieved by collecting human feedback [41, 72] on pairs of model-generated outputs and learning a function that predicts which output is preferred. Formally, given an input x and two outputs $y^{+}$ (preferred) and $y^{-}$ (dis-preferred), the model aims to satisfy:
@@
where $\pi_{\theta}(y|x)$ is the model’s probability of generating output $y$ given input $x$ with parameters $\theta$.
@@ 3 Temporal Preference Optimization
Unlike approaches that align image-LMM’s outputs with human preferences, our work focuses on aligning model outputs with intrinsic temporal preferences. Specifically, given a video input, the model is expected to produce responses that are better aligned with temporally grounded video content. To address this challenge, as illustrated in Figure 1, we introduce Temporal Preference Optimization (TPO), a framework designed to encourage the model to prioritize outputs that are more temporally grounded. To minimize reliance on extensive human annotations or advanced teacher models, we develop a scalable pipeline for generating temporal preference data by bootstrapping the model’s existing capabilities (Sec. 3.1).
Using this pipeline, we automatically generate paired datasets consisting of temporally preferred and dis-preferred examples. These pairs are then utilized for preference optimization (Sec. 3.5), enabling the model to learn to differentiate and favor temporally grounded outputs while preserving the capabilities of the original pre-trained model. Using this self-training approach, TPO effectively enhances temporal reasoning while maintaining scalability and practicality.
3.1 Temporal Preference Data Curation Overview
Formally, a temporal preference dataset $\mathcal{D}$ is constructed as a collection of tuples $(V,q,r^{+},r^{-})$, where $V$ denotes a video, $q$ represents a query, $r^{+}$ is the preferred temporally grounded response, and $r^{-}$ is the dis-preferred response. In this context, $r^{+}$ is designed to be more aligned with the temporal context of the video-query pair $(V,q)$ than $r^{-}$.
What constitutes good temporal grounding for video-LMMs? Effective temporal grounding in videos can be intuitively understood through two key scenarios: 1) Localized Temporal Grounding (Sec. 3.2): Given a question and a long video, the answer often pertains primarily to a small segment of the video. A video-LMM with strong temporal grounding capabilities should be able to localize and focus on this short duration, generating an accurate response without being distracted by irrelevant parts of the video. 2) Comprehensive Temporal Grounding (Sec. 3.3): When a question involves multiple keyframes across a video, the video-LMM is expected to identify and ground all relevant keyframes, ensuring no crucial information is overlooked.
Motivated by these two scenarios, we propose a multi-granularity temporal preference data generation pipeline that systematically captures these aspects of temporal grounding. This pipeline is designed to create temporal preference data that emphasizes both localized and comprehensive temporal grounding, enabling video-LMMs to excel across varying levels of temporal complexity.
3.2 Localized Temporal Preference Data
Query Generation
The queries are specifically crafted to focus on a subsegment of the video, ensuring they cannot be accurately answered using only general knowledge of the entire video. To generate such queries, we first sample a subsegment from the full video. Next, we sparsely sample $N$ frames from the selected subsegment and use an image-based vision-language model (CogVLM2 [20]) to generate captions for each frame. These captions serve as the basis for creating targeted questions. To ensure diversity and relevance, we design multiple question types and feed a structured question-generation prompt, along with the generated captions, into a LLM (GPT-4o-mini). This process produces a set of candidate questions specifically tailored to the sampled subsegment, facilitating the creation of localized temporal preference data.
Preferred Response Generation
@@ 3.2 Localized Temporal Preference Data
The preferred response in the curated dataset is expected to be of high quality and closely grounded in the corresponding temporal content. Current video-LMMs exhibit strong performance in understanding short video clips, making them well-suited to generate reliable responses when the video clip is both concise and directly relevant to the given query. To generate the preferred response, we provide the model with the same video segment used to curate the query, along with the query itself. By ensuring that the video clip is both brief and highly relevant to the query, we create conditions that maximize the likelihood of the model producing a high-quality, temporally grounded response. This process ensures that the preferred response aligns with the ideal characteristics for localized temporal grounding in video-LMMs.
Dis-Preferred Response Generation
The dis-preferred response in the preference dataset reflects the type of output the model aims to discourage. Specifically, it highlights responses where the model fails to localize relevant information in the video, thereby exposing shortcomings in temporal grounding. To generate the dis-preferred response, we clip out the query-related subsegment of the video and provide the model with the remaining portion of the video along with the query. Unlike the preferred response, where the model receives only the specific segment used to curate the query, this setup simulates an extreme scenario where the model entirely misses the relevant content. Consequently, the model must rely solely on the global context, high-possible producing responses with errors in temporal reasoning—precisely the kind of outputs our framework seeks to mitigate. This process ensures a clear contrast between preferred and dis-preferred responses, guiding the model to improve its temporal grounding capabilities.
3.3 Comprehensive Temporal Preference Data
Query Generation
Similar to the process used for localized temporal query generation, we employ a pre-trained image-LMM to generate captions. However, in this case, $N$ frames are sparsely sampled from the entire video to ensure broader coverage of the video content. These captions are then input into an LLM, which generates queries designed to capture comprehensive temporal relationships across multiple frames.
Preferred Response Generation
To generate the preferred response for the query in comprehensive TPO, we input the same sampled $N$ frames used during query generation into the video-LMM, ensuring an unaltered and exhaustive source of visual information. $N$ is set to a maximum of 32, striking a balance between information richness and generation precision. This complete visual input allows the video-LMM to effectively identify salient keyframes and perform reasoning across the entire temporal sequence of the video, producing comprehensive and accurate responses.
Dis-Preferred Response Generation
For the comprehensive task, we divide the video into subsegments and uniformly drop specific subsegments to create a modified version of the video with incomplete information. This truncated video, along with the comprehensive query, is then fed into the video-LMM to generate dis-preferred responses. Since the model lacks access to the full information relevant to the query, it often produces suboptimal responses based on the partial content. This approach simulates scenarios where the video-LMM fails to ground all critical video content necessary to answer comprehensive questions.
3.4 Preference Data Post Filtering
@@
After generating the queries and preference data pairs, some incorrect pairs may exist, such as cases where the dis-preferred response is equal to or better than the preferred response, or where the preferred response is incorrect in relation to the given query. To address this, we implement a post-filtering pipeline to enhance data quality and reduce noise. Using the LLM (GPT-4o-mini), we provide the captions along with their corresponding queries and preference data pairs, instructing the model to filter out pairs that meet these conditions. This post-filtering step allows us to eliminate corner cases that could introduce noise into the preference dataset, resulting in a more refined and higher-quality dataset that better supports the optimization process.
@@ 3.5 Training Objective
The generated preference dataset is subsequently leveraged to optimize the video-LMM’s temporal preferences using Direct Preference Optimization (DPO) [44], chosen for its flexibility and stability in handling preference-based learning tasks.
@@ 4.1 Experimental Settings
We evaluate TPO and baselines on three widely recognized benchmarks in multimodal video understanding, with particular attention to long-form video comprehension.
@@
Model LongVideoBench MLVU (M-avg) Video-MME
Short Medium Long Average
LongVA-7B [65] 51.3 58.8 61.1/61.6 50.4/53.6 46.2/47.6 52.6/54.3
@@
  + Hound-DPO [66, 65] 52.8 59.1 62.2/65.8 52.4/54.8 46.1/46.3 53.6/55.6
LongVA-TPO (ours) 54.2 61.7 63.1/66.6 54.8/55.3 47.4/47.9 55.1/56.6
Table 1: Results of LongVA-TPO on LongVideoBench [55], MLVU [70] and Video-MME [15] benchmarks compared to 3 baseline methods mentioned in 4.2. The Video-MME results are presented in the format “w/o subs / w/ subs”. The results for LongVA and LongVA+Hound-DPO are based on publicly available checkpoints, while the other results are evaluated using our trained model.
Model Size LongVideoBench MLVU Video-MME
(M-avg) Short Medium Long Average
GPT-4o [2] - 66.7 64.6 80.0/82.8 70.3/76.6 65.3/72.1 71.9/77.2
Video-LLaVA [31] 7B 39.1 47.3 45.3/46.1 38.0/40.7 36.2/38.1 39.9/41.6
LLaVA-1.5 [33] 7B 40.3 - - - - -
PLLaVA [57] 7B 40.2 - - - - -
Qwen-VL-Max [5] - - 42.2 55.8/57.6 49.2/48.9 48.9/47.0 51.3/51.2
ShareGPT4Video [7] 8B 39.7 46.4 48.3/53.6 36.3/39.3 35.0/37.9 39.9/43.6
InternVL-Chat-V1.5 [11] 20B 51.2 50.4 50.7/52.4 60.2/61.7 46.4/49.1 45.6/46.6
VideoChat2 [28] 7B 39.3 47.9 48.3/52.8 37.0/39.4 33.2/39.2 39.5/43.8
LongLLaVA [61] 7B - 56.3 61.9/66.2 51.4/54.7 45.4/50.3 52.9/57.1
Video-CCAM [14] 14B - 63.1 62.2/66.0 50.6/56.3 46.7/49.9 53.2/57.4
NVILA [37] 7B 57.7 70.1 75.7/77.6 62.2/69.0 54.8/63.3 64.2/70.0
Qwen2-VL [52] 7B 55.6 - - - - 63.3/69.0
Apollo [74] 7B 58.5 70.9 - - - 61.3/63.3
LongVA-7B [65] 7B 51.3 58.8 61.1/61.6 50.4/53.6 46.2/47.6 52.6/54.3
LLaVA-Video-7B [68] 7B 58.2 70.8 - - - 63.3/69.7
LongVA-TPO (ours) 7B 54.2 61.7 63.1/66.6 54.8/55.3 47.4/47.9 55.1/56.6
LLaVA-Video-TPO (ours) 7B 60.1 71.1 76.8/78.7 64.6/69.4 55.4/66.4 65.6/71.5
Table 2: Results on LongVideoBench [55], MLVU [70] and Video-MME [15] compared with state-of-the-art models. The Video-MME results are presented in the format “w/o subs / w/ subs”.
@@
The model is trained on 8 Nvidia A100 80GB GPUs, with a batch size of 1. For the preference optimization on LongVA, we set the KL-divergence weight $\beta=0.3$ and the SFT loss weight $\alpha=0.5$, while for LLaVA-Video, we set the KL-divergence weight $\beta=0.2$ and the SFT loss weight $\alpha=1$. We train the model on our curated data for 1 epoch. It takes about 4 hours for TPO to perform on LongVA-7B with a learning rate of $4e^{-6}$ and also about 4 hours for LLaVA-Video-7B with a learning rate of $3e^{-7}$. During data preparation, we employ the GPT-4o-mini language model (text-only input) for question curation and post-filtering. This choice balances cost-effectiveness with efficiency, facilitating a streamlined and scalable data processing workflow.

### v3

# Temporal Preference Optimization of Large Multimodal Models

@@ Abstract
Abstract: Despite recent advancements in video Large Multimodal Models (video-LMMs), accurate temporal grounding remains a key challenge. In this work, we introduce Temporal Preference Optimization (TPO)—a post-training framework that unlocks superior temporal reasoning in video-LMMs without requiring human annotations. TPO enables preference modeling by manipulating video inputs to generate contrastive responses, ensuring that preferred responses are more temporally grounded than dis-preferred ones. Through preference learning, TPO enhances the model’s capability to distinguish and localize events with better temporal reasoning. Extensive experiments on LongVideoBench, MLVU, and Video-MME demonstrate that TPO significantly improves temporal grounding across multiple video-LMMs. Notably, LLaVA-Video-TPO achieves state-of-the-art performance among 7B models on Video-MME, establishing TPO as a scalable and effective solution for advancing temporal understanding in video analysis.
Equal Contributions.
Figure 1: Temporal Preference Optimization (TPO). TPO is a post-training algorithm designed to enhance temporal comprehension in video-LMMs. In TPO, (a) preference data is generated by prompting the video-LMM with both well-grounded and manipulated (irrelevant or incomplete) video clips to collect contrastive response pairs. (b) These pairs undergo an LLM-based post-filtering process to remove noisy or misaligned samples. The generated preference data is then used in the preference optimization process, which improves the model’s temporal reasoning by prioritizing preferred responses, ultimately enhancing overall video understanding.
[…]
@@ 1 Introduction
Recent advancements in video Large Multimodal Models (video-LMMs) [52, 2, 45] have marked a significant step forward for generalizable video understanding. While image-based LMMs [19, 5, 39] primarily focus on spatial reasoning, video-LMMs face the additional complexity of modeling temporal dependencies—a critical aspect for capturing the dynamic nature of video content.
Most existing video-LMMs acquire temporal grounding implicitly during supervised finetuning by leveraging weak correspondences between input videos and textual responses [6, 67]. While some responses may reference specific segments of a video, they lack explicit temporal alignment under next-token prediction training, limiting the model’s ability to learn precise temporal grounding. Recently, alternative approaches [46, 7, 29, 20, 51] have emerged that incorporate explicit temporal annotations into training, enriching textual responses with structured temporal information as supervision. However, these methods rely on additional temporal annotations, which are costly to obtain and scale to large training datasets.
In this work, we introduce Temporal Preference Optimization (TPO), a post-training framework designed to enhance the temporal grounding capabilities of video-LMMs. TPO systematically refines the pretrained video-LMMs’ ability to distinguish temporally relevant content by leveraging contrastive responses from manipulated video inputs. As shown in Fig. 1, TPO first prompts a video-LMM with the same question on both the original and the corrupted video. Questions are formulated based on a set of video frames, with preferred responses generated using these frames. In contrast, dis-preferred responses are produced using the same question but paired with irrelevant or incomplete frame sets. This pipeline ensures that preferred responses contain richer and more temporally relevant information than dis-preferred ones, thereby establishing a clear preference hierarchy. By dynamically manipulating video inputs based on the query, TPO automatically injects temporal preferences into the preference dataset through simple input transformations. To further refine the dataset, a post-filtering process is applied to remove imperfect samples caused by errors from the pretrained video-LMMs and ambiguous preference data. The resulting preference dataset is then used to optimize the model’s temporal grounding capabilities via Direct Preference Optimization (DPO) [44], chosen for its flexibility and stability. This structured pipeline enables TPO to enhance the temporal reasoning capabilities of the base video-LMM using a curated post-training dataset, while preserving its pre-trained knowledge. This makes TPO a scalable and robust solution for advancing video understanding tasks.
@@
We conducted extensive experiments on three challenging multimodal video understanding benchmarks, and the results clearly demonstrate that TPO significantly enhances the temporal grounding capabilities of video-LMMs. Specifically, TPO achieves performance gains of 2.9% on LongVideoBench [55], 3.1% on MLVU [70], and 2.5% on Video-MME [14], when applied to the strong base model LongVA-7B [65]. Furthermore, even when integrated with the state-of-the-art large-scale pretrained video-LMM, LLaVA-Video, TPO still delivers a 2.3% improvement, establishing LLaVA-Video-TPO as the best-performing 7B model on the Video-MME benchmark.
@@ 2 Preliminaries
Preference learning [41, 49, 72] focuses on modeling human preferences to align model behavior with user expectations. In LLMs and image-LMMs, this involves training models to generate responses favored by users. This is typically achieved by collecting human feedback on pairs of model-generated outputs and learning a function that predicts which output is preferred. Formally, given an input x and two outputs $y^{+}$ (preferred) and $y^{-}$ (dispreferred), the model aims to satisfy:
@@
where $\pi_{\theta}(y|x)$ is the model’s probability of generating output $y$ given input $x$ with parameters $\theta$. In the context of video-LMMs, a preference dataset $\mathcal{D}$ is constructed as a collection of tuples $(V,q,r^{+},r^{-})$, where $V$ denotes a video, $q$ represents a query, $r^{+}$ is the preferred temporally grounded response, and $r^{-}$ is the dispreferred response.
@@ 3 Temporal Preference Optimization
While prior works focus primarily on aligning LLM outputs with human preferences, our approach uniquely aligns model outputs with intrinsic temporal preferences in videos. To achieve this, we propose Temporal Preference Optimization (TPO), illustrated in Fig. 2, which significantly enhances the video reasoning capabilities of video-LMMs. TPO systematically incorporates temporal grounding into the optimization process through creating preference pairs from the contrast between meticulously manipulated video inputs (Sec. 3.1). To further enhance the quality of these preference data, we introduce a rule-based post-filtering step (Sec. 3.2). Finally, Direct Preference Optimization (Sec. 3.3) is leveraged to optimize the model towards temporally preferred outputs without compromising its original pretrained capabilities.
3.1 Temporal Preference Modeling
Query Generation.
Given a video $\bf{V}$, we first sample a segment containing a set of frames $S_{a}$, which may be a subset of the video or the entire sequence of frames. To generate descriptive context, we employ an image-based vision-language model (CogVLM2 [19]) to generate captions for each frame in $S_{a}$. These captions serve as the foundation for constructing targeted questions. To ensure diversity and relevance, we design multiple question types and use a structured question-generation prompt, as shown in Fig. 8 (Appendix), incorporating the generated captions. This prompt is then processed by a large language model (GPT-4o-mini) to produce a set of candidate questions specifically tailored to the sampled video frames, resulting in a set of questions $S_{q}$. This approach ensures that the generated questions are contextually relevant that allows precise control over subsequent response generation.
Figure 2: In TPO, we introduce two approaches for temporal preference data generation. Preferred responses are generated using questions and their corresponding frames for strong temporal grounding. For dis-preferred responses, we introduce: (a) Generation with Irrelevant Information, where all relevant frames are excluded. (b) Generation with Incomplete Information, where only a partial subset of relevant frames is used. These manipulated clips create contrastive response pairs, highlighting differences between well-grounded and manipulated video content. This contrast serves as a learning signal to enhance the model’s temporal reasoning.
Preferred Response Generation.
@@ 3.1 Temporal Preference Modeling
Preferred responses in the curated dataset are expected to be strongly grounded in the corresponding temporal content. To achieve this, we use the question set $S_{q}$ along with their corresponding frame set $S_{a}$ as input to the video-LMM. By ensuring that the provided video frames are highly relevant to the query, we create conditions that maximize the likelihood of the model generating a high-quality, temporally grounded response. This process guarantees that the preferred responses align with the ideal characteristics for effective temporal grounding in video-LMMs.
Dis-Preferred Response Generation.
The dis-preferred responses in the preference dataset represent the type of outputs the model aims to discourage, specifically those where it fails to localize relevant information in the video. These responses expose shortcomings in temporal reasoning by highlighting cases where the model struggles to align its predictions with the actual video content. To generate these dis-preferred responses, we manipulate the video inputs to simulate imperfect temporal grounding. As illustrated in Fig. 2, we introduce two strategies for constructing the input frame set $S_{b}$ used in dis-preferred response generation:
(a) Generation with Irrelevant Information: To simulate an extreme failure case where the model misses all relevant frames, we construct $S_{b}$ by excluding the relevant frame set $S_{a}$ and instead sampling from the remaining frames of the video. This ensures that $S_{b}$ contains only irrelevant content, forcing the model to generate a response based on unrelated visual information.
(b) Generation with Incomplete Information: To mimic scenarios where the model has access to only partial relevant information, we construct $S_{b}$ by randomly sampling a subset of $S_{a}$. This setup introduces gaps in the temporal context, making it harder for the model to fully comprehend the event described in the query.
Unlike preferred responses, which are grounded in fully relevant video segments, these manipulated setups significantly increase ambiguity and noise by partially or completely omitting critical visual content. As a result, the model is forced to rely on incomplete or misleading information, making temporal reasoning errors and hallucinations more likely. This intentional contrast between preferred and dis-preferred responses serves as a strong learning signal, helping refine the model’s ability to distinguish and accurately localize events in time, ultimately enhancing its temporal reasoning capabilities.
3.2 LLM-based Post-Filtering
Although we design the preferred responses to be of higher quality than the dis-preferred responses, this distinction is not always guaranteed due to the imperfect video understanding capabilities of the base video-LMMs. In some cases, errors in response generation may lead to misaligned preference pairs, where the preferred response contains noise or the dis-preferred response is of better quality than expected.
@@ 3.2 LLM-based Post-Filtering
To enhance data quality and reduce noise, we introduce a post-filtering pipeline leveraging an LLM (GPT-4o-mini). Specifically, we provide the model with the key frame captions of $S_{a}$, along with their corresponding queries and preference data pairs, and instruct it to filter out samples that meet predefined criteria (detailed prompts are shown in Fig. 9 in the Appendix). The filtering rules target cases where: 1) The dis-preferred response is of higher quality than the preferred response. 2) The preferred response is factually incorrect or misaligned with the video content. 3) The query itself is ambiguous, making preference ranking unreliable. By incorporating this post-filtering step, we effectively eliminate problematic cases that could introduce noise into training, resulting in a refined, higher-quality dataset that better supports effective model optimization and improves temporal grounding performance.
@@ 3.3 Training Objective
The generated preference dataset is leveraged to optimize the temporal grounding capabilities of video-LMMs using Direct Preference Optimization (DPO) [44], selected for its robustness and effectiveness in preference-based learning.
@@ 4.1 Experimental Settings
We evaluate TPO and baselines on three widely recognized benchmarks in multimodal video understanding.
@@
Model Size LongVideo MLVU Video-MME
Bench (M-avg) Short Medium Long Average
GPT-4o [2] - 66.7 64.6 80.082.8 70.376.6 65.372.1 71.977.2
Video-LLaVA [31] 7B 39.1 47.3 45.346.1 38.040.7 36.238.1 39.941.6
LLaVA-1.5 [33] 7B 40.3 - - - - -
PLLaVA [57] 7B 40.2 - - - - -
Qwen-VL-Max [5] - - 42.2 55.857.6 49.248.9 48.947.0 51.351.2
ShareGPT4Video [6] 8B 39.7 46.4 48.353.6 36.339.3 35.037.9 39.943.6
InternVL-Chat-V1.5 [9] 20B 51.2 50.4 50.752.4 60.261.7 46.449.1 45.646.6
VideoChat2 [28] 7B 39.3 47.9 48.352.8 37.039.4 33.239.2 39.543.8
LongLLaVA [61] 7B - 56.3 61.966.2 51.454.7 45.450.3 52.957.1
Video-CCAM [13] 14B - 63.1 62.266.0 50.656.3 46.749.9 53.257.4
NVILA [36] 7B 57.7 70.1 75.777.6 62.269.0 54.863.3 64.270.0
Qwen2-VL [52] 7B 55.6 - - - - 63.369.0
Apollo [74] 7B 58.5 70.9 - - - 61.363.3
LongVA-7B [65] 7B 51.3 58.8 61.161.6 50.453.6 46.247.6 52.654.3
LLaVA-Video-7B [68] 7B 58.2 70.8 - - - 63.369.7
LongVA-TPO (ours) 7B 54.2 61.7 63.166.6 54.855.3 47.447.9 55.156.6
[…]
Table 1: Results on LongVideoBench [55], MLVU [70] and Video-MME [14] compared with state-of-the-art models. The Video-MME results are presented in the format w/o subsw/subs.
Model LongVideoBench MLVU Video-MME
(M-avg) Short Medium Long Average
LongVA-7B [65] 51.3 58.8 61.161.6 50.453.6 46.247.6 52.654.3
@@
  + Hound-DPO  [66, 65] 52.8 59.1 62.265.8 52.454.8 46.146.3 53.655.6
  + Hound-DPO  [66, 65] 52.6 59.3 63.165.9 50.854.7 47.247.0 53.755.9
[…]
Table 2: Results of LongVA-TPO on LongVideoBench [55], MLVU [70] and Video-MME [14] benchmarks compared to 3 baseline methods mentioned in 4.2. The Video-MME results are presented in the format “w/o subs / w/ subs". The results for LongVA and LongVA+Hound-DPO are based on publicly available checkpoints, and for LongVA+Hound-DPO are based on our implementation on our collected video datasets, while the other results are evaluated using our trained model.
@@
The model is trained on 8 Nvidia A100 80GB GPUs, with a batch size of 64. For the preference optimization on LongVA, we set the KL-divergence weight $\beta=0.3$ and the SFT loss weight $\alpha=0.5$, while for LLaVA-Video, we set the KL-divergence weight $\beta=0.2$ and the SFT loss weight $\alpha=1$. We train the model on our curated data for 1 epoch. It takes about 4 hours for TPO to perform on LongVA-7B with a learning rate of $4e^{-6}$ and also about 4 hours for LLaVA-Video-7B with a learning rate of $3e^{-7}$. During data preparation, we employ the GPT-4o-mini language model (text-only input) for question curation and post-filtering. This choice balances cost-effectiveness with efficiency, facilitating a streamlined and scalable data processing workflow.